NeilArmweak:

The terminal side of angle in standard position intersects the unit circle at each point P . Find cos and sin . P(0, –1) A: cos θ = –1, sin θ = 0 B:: cos θ = 0, sin θ = C:: cos θ = 0, sin θ = 1 D: cos θ = 0, sin θ = –1

NeilArmweak:

My thoughts on this is that the answer might be D

Extrinix:

Is there a image of this?

NeilArmweak:

no they didn't provide it

Extrinix:

Okay.

NeilArmweak:

I think it's D though?

Extrinix:

\(cos(\theta) = x-coordinate\) \(sin(\theta)=y-coordinate\) Using this, which one do you think it is?

NeilArmweak:

D

NeilArmweak:

wait x minus coordinate?

surjithayer:

 \[\cos \theta=\cos 270=\cos (180+90)=-\cos 90=0\] \[\sin \theta =\sin (270)=\sin(180+90)=-\sin 90=-1\] where theta is the angle which the terminal line makes with initial line.